# Scaling Up Liquid-Resistance Liquid-Capacitance Networks for Efficient Sequence Modeling

### v2

@@ 3.2 Comparison to Linear State Space Models
State-of-the-art time-invariant LSSMs typically take the following general form:
@@
The main differences between LrcSSM and time-invariant LSSMs are the following:
  • There is no non-linearity in the recurrent and input update ($\mathbf{A}$ and $\mathbf{B}$, respectively) in time-invariant LSSMs, which allows them to be parallelized over the time dimension. Here, we investigate non-linear recurrent update and non-linear input update too.
  • There are two key aspects of the matrices that state-of-the-art LSSMs usually follow:
@@
      First, matrix $\mathbf{A}$ is generally time-invariant (constant), although recent work has introduced an input-dependent variant $\mathbf{A(u)}$ [14, 9]. In our model however, this matrix is both state-and-input dependent, $\mathbf{A}(\mathbf{x},\mathbf{u})$. Second, instead of using a traditional $\mathbf{B}$ matrix that is simply multiplied by the input $\mathbf{u}$, we adopt the form $\mathbf{b}(\mathbf{x},\mathbf{u})$, allowing the input to have a more embedded influence on the state update.
@@
      Modern LSSMs typically require a special initialization, such as diagonal plus-low rank parameterization of the transition matrix of the LSSMs via higher-order polynomial projection (HiPPO) matrix [10] or only diagonal transition matrices with specific parameterization [11, 28]. In our case, we calculate the entries of $\mathbf{A}$ and $\mathbf{b}$ from biology-grounded equations of (13) and (14).
@@ 3.3 Comparison to Liquid-Resistance Liquid-Capacitance Networks (LRCs)
  • Learning in LRCs, like in traditional NSSMs (or nonlinerar RNN models), is inherently sequential. In contrast, we aim for an efficient, parallelizable version in LrcSSMs.
@@
  • While the original LRCs use a single computation layer (a single computation block), we have restructured the LRC architecture into a block-wise design in LrcSSMs, similar to the LSSM-styled models such as LRU and S5. This design is illustrated in Figure 1.
@@ 3.4 Theoretical Insights
The LrcSSM architecture enjoys three important theoretical properties. Firstly, by forcing the state-transition matrix of LrcSSMs to be diagonal and learned at every time step, the full sequence can be solved in parallel with a single prefix-scan, giving $\mathcal{O}(TD)$ time and memory and only $\mathcal{O}(\log T)$ sequential depth, where $T$ is the input-sequence length, and $D$ is the state dimension.
@@
The full proof of all these properties is given in Appendix A, due to obvious space limitations. In particular we provide all details about LrcSSMs stability in A.1 and scalability in A.2.
@@ 4 Related work
Linear Structural State-Space Models (LSSMs).
Since the introduction of S4 [12], LSSMs rapidly evolved in sequence modeling. S4 used FFT to efficiently solve linear recurrences, and inspired several variants, including S5 [34], which replaced FFT with parallel scans. Liquid-S4 [14] introduced input-dependent state-transition matrices, moving beyond the static structure but relied on FFT. Recent work, such as S6 and Mamba [9], adapted the concept of input-dependency and continued to push LSSMs to more efficient computation with a hardware-aware parallel algorithm.
Parallelizing Non-linear State Space Models (NSSMs).
While traditional nonlinear RNNs have been favored for their memory efficiency, their major limitation lies in the lack of parallelizability over the sequence length. This has led to the development of parallelizable alternatives, such as [25, 28, 27, 8]. One notable example is the Linear Recurrent Unit (LRU) [28], which uses complex diagonal state-transition matrices with stable exponential parameterization, achieving comparable performance with LSSMs. While LRUs argue that linear recurrence is sufficient, in this work we show that incorporating non-linearity in the transition dynamics can offer significant advantages. Importantly, these approaches achieve parallelism through entirely new architectures, without addressing how to parallelize existing NSSMs. Techniques like DEER [24] and ELK [7] fill this gap by enabling parallel training and inference for arbitrary non-linear recurrent models.
@@ 5 Experiments
We compare LrcSSMs against nine models representing the state of the art for a range of long-sequence tasks. These include the Neural Controlled Differential Equations (NCDE) [20], Neural Rough Differential Equations (NRDE) [26] and Log-NCDE [35], Linear Recurrent Unit (LRU) [28], S5 [34], MAMBA [9], S6 [9], Linear Oscillatory State-Space models with implicit-explicit time integration (LinOSS-IMEX [32]) and with implicit time integration (LINOSS-IM [32]).
We follow the same classification evaluation benchmark proposed in [35] and then used by [32]. These tasks are part of the UEA Multivariate Time Series Classification Archive (UEA-MTSCA). All of these datasets consist of biologically or physiologically grounded time-series data, derived from real-world measurements of dynamic systems, which can be human, animal, or chemical. They capture continuous temporal signals such as neural activity, bodily movements, or spectroscopic readings, making them very well-suited for benchmarking models that need to learn complex temporal dependencies. We followed the exact same hyperparameter-tuning protocol, using a grid search over the validation accuracy. More details on these experiments are given in the Appendix B.2. After fixing the hyperparameters, we compare the average test set accuracy over five different random splits of the data. As we are reporting the results of the other models from Rusch et. al [32], we also used the exact same seeds for the dataset splitting as well. When presenting the results, we highlight the top three performing models.
@@
In Table 1, we report results on datasets with sequence lengths shorter than 1,500 elements. These datasets include the Heartbeat dataset [6], which contains heart sound recordings, as well as SelfRegulationSCP1 and SelfRegulationSCP2 [1], which include data on cortical potentials. We report the test accuracy results. We found that our LrcSSM model performed average on these tasks, and we suspect that they lack interesing input correlations.
Table 1: Test accuracy comparison of different models across relatively short-horizon datasets (<1,500). The performance of the models marked by † is reported from [32]. The same hyperparameter tuning protocol and dataset splitting over the same 5 seeds were used.
Heartbeat SelfRegulationSCP1 SelfRegulationSCP2
Sequence length 405 896 1,152
Input size 61 6 7
#Classes 2 2 2
NRDE 73.9 $\pm$ 2.6 76.7 $\pm$ 5.6 48.1 $\pm$ 11.4
NCDE 68.1 $\pm$ 5.8 80.0 $\pm$ 2.0 49.1 $\pm$ 6.2
Log-NCDE 74.2 $\pm$ 2.0 82.1 $\pm$ 1.4 54.0 $\pm$ 2.6
LRU 78.1 $\pm$ 7.6  84.5 $\pm$ 4.6 47.4 $\pm$ 4.0
S5 73.9 $\pm$ 3.1 87.1 $\pm$ 2.1 55.1 $\pm$ 3.3
Mamba 76.2 $\pm$ 3.8 80.7 $\pm$ 1.4 48.2 $\pm$ 3.9
S6 76.5 $\pm$ 8.3 82.8 $\pm$ 2.7 49.9 $\pm$ 9.4
LinOSS-IMEX 75.5 $\pm$ 4.3 87.5 $\pm$ 4.0 58.9 $\pm$ 8.1
LinOSS-IM 75.8 $\pm$ 3.7 87.8 $\pm$ 2.6 58.2 $\pm$ 6.9
LrcSSM (Ours) 72.7 $\pm$ 5.7 85.2 $\pm$ 2.1 53.9 $\pm$ 7.2
@@
Next, we focus on the tasks that require learning long-range interactions, especially those with a sequence length above 1,500, up to 18,000. These include the EthanolConcentration dataset [22], which contains spectroscopic recordings of solutions, the MotorImagery dataset, which captures data from the motor cortex, and the EigenWorms [38] dataset of postural dynamics of the worm C. elegans. As shown in Table 2, our LrcSSM model outperforms all other state-of-the-art methods on the EthanolConcentration task and achieves second-best performance on the MotorImagery and EigenWorms datasets. We believe that EthanolConcentration contains interesting input correlations, which LrcSSMs can capture through its state dependence.
Table 2: Test accuracy comparison of different models across long-horizon datasets (>1,500). The performance of the models marked by † is reported from [32]. Results are averaged over 5 seeds.
EthanolConcentration MotorImagery EigenWorms
Sequence length 1,751 3,000 17,984
Input size 2 63 6
#Classes 4 2 5
NRDE 31.4 $\pm$ 4.5 54.0 $\pm$ 7.8 77.2 $\pm$ 7.1
NCDE 22.0 $\pm$ 1.0 51.6 $\pm$ 6.2 62.2 $\pm$ 2.2
Log-NCDE 35.9 $\pm$ 6.1 57.2 $\pm$ 5.6 82.8 $\pm$ 2.7
LRU 23.8 $\pm$ 2.8 51.9 $\pm$ 8.6 85.0 $\pm$ 6.2
S5 25.6 $\pm$ 3.5 53.0 $\pm$ 3.9 83.9 $\pm$ 4.1
Mamba 27.9 $\pm$ 4.5 47.7 $\pm$ 4.5 70.9 $\pm$ 15.8
S6 26.4 $\pm$ 6.4 51.3 $\pm$ 4.7 85.0 $\pm$ 16.1
LinOSS-IMEX 29.9 $\pm$ 1.0 57.9 $\pm$ 5.3 80.0 $\pm$ 2.7
LinOSS-IM 29.9 $\pm$ 0.6 60.0 $\pm$ 7.5 95.0 $\pm$ 4.4
LrcSSM (Ours) 36.9 $\pm$ 5.3 58.6 $\pm$ 3.1 90.6 $\pm$ 1.4
@@
In Table 3, we report the average accuracy across all six datasets considered from the UEA-MTSCA archive. LrcSSM achieved an accuracy of $66.3\%$, placing it at the forefront alongside the LinOSS-IM model, outperforming all other state-of-the-art models, including LRU, S5, S6, Mamba, and LinOSS-IMEX. The implicit integration scheme of the LinOSS-IM model, seems to have played an important role, and we plan to investigate a similar integration scheme for LrcSSM, too. Our current scheme is just a simple explicit Euler.
Table 3: Average test accuracy (%) across all datasets. As before, the performance of the models marked by † is reported from [32]. Results are averaged over 5 seeds.
Average Test Accuracy
NRDE 60.2 $\pm$ 17.1
NCDE 55.5 $\pm$ 18.2
Log-NCDE 64.4 $\pm$ 16.9
LRU 61.8 $\pm$ 22.6
S5 63.1 $\pm$ 21.2
Mamba 58.6 $\pm$ 18.8
S6 62.0 $\pm$ 21.2
LinOSS-IMEX 65.0 $\pm$ 19.0
LinOSS-IM 67.8 $\pm$ 21.6
LrcSSM (Ours) 66.3 $\pm$ 18.6
@@ 6 Discussion
Our experimental evaluations show that the LrcSSM model performs moderately well on short-horizon datasets, as seen in Table 1, while demonstrating highly competitive performance on datasets with long input sequences, as shown in Table 2. In those long-sequence tasks, LrcSSMs outperform LRUs, Mamba, and S6, and also achieve better average performance across all datasets, as presented in Table 3.
The only model that LrcSSMs generally does not outperform is the LinOSS-IM model, except on the EthanolConcentration dataset (for both LinOSS-IMEX and LinOSS-IM versions), and on MotorImagery and EigenWorms (in the case of LinOSS-IMEX). This may be attributed to the fact that LinOSS is based on forced linear second-order ODEs, whereas LrcSSMs are built upon LRCs, which are nonlinear first-order ODEs. Another possible reason lies in the integration technique: while we were able to outperform the implicit-explicit (IMEX) integration scheme, we did not surpass the fully implicit one (IM) in average test accuracy. This suggests that more sophisticated integration schemes for LrcSSMs (which currently use explicit Euler) may be worth investigating.
@@
In this paper, we focused on the biologically inspired non-linear LRC model, and demonstrated how this model can be made more efficient for long-sequence modeling, by redesigning its underlying state-transition matrix $\mathbf{A}$ and its input-transition vector $\mathbf{b}$, such that the resulting Jacobian is a diagonal matrix, for the state-update iterations. This matrix can then be directly used in the parallelizable ELK method, which gives an exact ELK update, and not an approximation. We believe this approach can also be applied to many other non-linear RNNs of interest.
@@
As pointed out in Section A.2, this is parallelized version holds a good promise towards efficient non-linear RNNs compared to sequential computation costs. Linear SSMs have also the same costs. However, we also have to take into account that LRCs solved by ELK need more Newton steps to converge at each iteration, which linear SSMs do not require. The number of iterations depends on the convergence of the state updates, which stops once the difference between the consecutive state updates gets below a defined threshold (see Line 4 of Algorithm 1).

### v6

@@ 3.2 Comparison to Linear State Space Models
State-of-the-art time-invariant linear SSMs typically take the following general form:
@@
The main differences between LrcSSM and time-invariant linear SSMs are the following:
  • There is no non-linearity in the recurrent state and input update ($\mathbf{A}$ and $\mathbf{B}$, respectively) in time-invariant linear SSMs, which allows them to be parallelized over the time dimension. Here, we investigate non-linear recurrent update and non-linear input update too.
  • There are two key aspects of the matrices that state-of-the-art linear SSMs usually follow:
@@
      First, matrix $\mathbf{A}$ is generally time-invariant (constant), although recent work has introduced an input-dependent variant $\mathbf{A(u)}$ [15, 10]. In our model however, this matrix is both state- and input-dependent, $\mathbf{A}(\mathbf{x},\mathbf{u})$. Second, instead of using a traditional $\mathbf{B}$ matrix that is simply multiplied by the input $\mathbf{u}$, we adopt the form $\mathbf{b}(\mathbf{x},\mathbf{u})$, allowing the input to have a more embedded influence on the state update.
@@
      Modern linear SSMs typically require a special initialization, such as diagonal plus low rank parameterization of the state matrix of the linear SSMs via higher-order polynomial projection (HiPPO) matrix [11] or only diagonal state matrices with specific parameterization [12, 32]. In our case, we calculate the entries of $\mathbf{A}$ and $\mathbf{b}$ from biology-grounded equations of (13) and (14).
@@ 3.3 Comparison to Liquid-Resistance Liquid-Capacitance Networks (LRCs)
  • Learning in LRCs, like in traditional non-linear RNN models, is inherently sequential. In contrast, we aim for an efficient, parallelizable version in LrcSSMs.
@@
  • While the original LRCs use a single computation layer (a single computation block), we have restructured the LRC architecture into a block-wise design in LrcSSMs, similar to the linear SSM-styled models such as LRU and S5. This design is illustrated in Figure 1.
@@ 3.4 Theoretical Insights
The LrcSSM architecture enjoys three important theoretical properties. Firstly, by forcing the state-transition matrix of LrcSSMs to be diagonal and learned at every time step, the full sequence can be solved in parallel, giving $\mathcal{O}(TD)$ time and memory and only $\mathcal{O}(\log T)$ sequential depth, where $T$ is the input-sequence length, and $D$ is the state dimension.
@@
The full proof of all these properties is given in Appendix A. In particular we provide details about LrcSSMs stability in A.1 and scalability in A.2.
@@ 4 Related Work
Linear Structural State-Space Models.
Since the introduction of S4 [13], linear SSMs rapidly evolved in sequence modeling. S4 used FFT to efficiently solve linear recurrences, and inspired several variants, including S5 [39], which replaced FFT with parallel scans. Liquid-S4 [15] introduced input-dependent state matrices, moving beyond the static structure but relied on FFT. Recent work, such as S6 and Mamba [10], adapted the concept of input-dependency and continued to push linear SSMs to more efficient computation with a hardware-aware parallel algorithm.
Parallelizing Non-linear Recurrent Neural Networks.
While traditional non-linear RNNs have been favored for their memory efficiency, their major limitation lies in the lack of parallelizability over the sequence length. This has led to the development of parallelizable alternatives, such as [28, 32, 31, 9]. One notable example is the Linear Recurrent Unit (LRU) [32], which uses complex diagonal recurrent state matrices with stable exponential parameterization, achieving comparable performance with SSMs. While LRUs argue that linear recurrence is sufficient, in this work we show that incorporating non-linearity in the transition dynamics can offer significant advantages. Importantly, these approaches achieve parallelism through entirely new architectures, without addressing how to parallelize existing non-linear RNNs. Techniques like DEER [27] and ELK [8] fill this gap by enabling parallel training and inference for arbitrary non-linear recurrent models.
@@ 5 Experiments
We follow the same classification evaluation benchmark proposed in [41] and then used by [37]. These tasks are part of the UEA Multivariate Time Series Classification Archive (UEA-MTSCA). All of these datasets consist of biologically or physiologically grounded time-series data, derived from real-world measurements of dynamic systems, which can be human, animal, or chemical. They capture continuous temporal signals such as neural activity, bodily movements, or spectroscopic readings, making them very well-suited for benchmarking models that need to learn complex temporal dependencies.
5.1 Short- and Long-Horizon Sequence Tasks
We compare LrcSSMs against eleven models representing the state of the art for a range of long-sequence tasks. These include the Neural Controlled Differential Equations (NCDE) [22], Neural Rough Differential Equations (NRDE) [30] and Log-NCDE [41], Linear Recurrent Unit (LRU) [32], S5 [39], MAMBA [10], S6 [10], Linear Oscillatory State-Space models with implicit-explicit time integration (LinOSS-IMEX [37]) and with implicit time integration (LINOSS-IM [37]), Transformers [40] and RFormers [29].
We followed the exact same hyperparameter-tuning protocol of [37], using a grid search over the validation accuracy. More details on these experiments are given in Appendix B.2. After fixing the hyperparameters, we compare the average test set accuracy over five different random splits of the dataset. As we are reporting the most of the model results from Rusch et. al [37], we also used the exact same seeds for the dataset splitting as well. When presenting the results, we highlight the top three performing models.
@@ 5.1 Short- and Long-Horizon Sequence Tasks
In the first block of Table 1, we report results on datasets with sequence lengths shorter than 1,500 elements. These datasets include the Heartbeat dataset (Heart) [7], which contains heart sound recordings, as well as SelfRegulationSCP1 (SCP1) and SelfRegulationSCP2 (SCP2) [1], which include data on cortical potentials. We report the test accuracy results. We found that our LrcSSM model performed average on these tasks, and we suspect that these datasets lack interesting input correlations.
@@
Next, we focus on the tasks that require learning long-range interactions, especially those with a sequence length above 1,500, up to 18,000. These include the EthanolConcentration dataset (Ethanol) [25], which contains spectroscopic recordings of solutions, the MotorImagery dataset (Motor) [24], which captures data from the motor cortex, and the EigenWorms (Worms) [44] dataset of postural dynamics of the worm C. elegans. As shown in the second block of Table 1, our LrcSSM model outperforms all other state-of-the-art SSM methods on the EthanolConcentration task and achieves second-best performance on the MotorImagery and EigenWorms datasets too. We believe that EthanolConcentration contains interesting input correlations, which LrcSSMs can capture through its input- and state-dependence.
Table 1: Test accuracy comparison of different models across relatively short-horizon datasets (<1,500) in the left block and long-horizon datasets (>1,500) in the middle block. Average test accuracy with standard error (%) across all datasets is in the rightmost block. The performance of the models marked by † is reported from [37], and those with ‡ from [29]. The same hyperparameter tuning protocol and dataset splitting over the same 5 seeds were used.
Heart SCP1 SCP2 Ethanol Motor Worms Average
Sequence length 405 896 1,152 1,751 3,000 17,984
Input size 61 6 7 2 63 6
#Classes 2 2 2 4 2 5
NRDE 73.9 $\pm$ 2.6 76.7 $\pm$ 5.6 48.1 $\pm$ 11.4 31.4 $\pm$ 4.5 54.0 $\pm$ 7.8 77.2 $\pm$ 7.1 60.2 $\pm$ 7.7
NCDE 68.1 $\pm$ 5.8 80.0 $\pm$ 2.0 49.1 $\pm$ 6.2 22.0 $\pm$ 1.0 51.6 $\pm$ 6.2 62.2 $\pm$ 2.2 55.5 $\pm$ 8.1
Log-NCDE 74.2 $\pm$ 2.0 82.1 $\pm$ 1.4 54.0 $\pm$ 2.6 35.9 $\pm$ 6.1 57.2 $\pm$ 5.6 82.8 $\pm$ 2.7 64.4 $\pm$ 7.6
LRU 78.1 $\pm$ 7.6 84.5 $\pm$ 4.6 47.4 $\pm$ 4.0 23.8 $\pm$ 2.8 51.9 $\pm$ 8.6 85.0 $\pm$ 6.2 61.8 $\pm$ 10.1
S5 73.9 $\pm$ 3.1 87.1 $\pm$ 2.1 55.1 $\pm$ 3.3 25.6 $\pm$ 3.5 53.0 $\pm$ 3.9 83.9 $\pm$ 4.1 63.1 $\pm$ 9.5
Mamba 76.2 $\pm$ 3.8 80.7 $\pm$ 1.4 48.2 $\pm$ 3.9 27.9 $\pm$ 4.5 47.7 $\pm$ 4.5 70.9 $\pm$ 15.8 58.6 $\pm$ 8.4
S6 76.5 $\pm$ 8.3 82.8 $\pm$ 2.7 49.9 $\pm$ 9.4 26.4 $\pm$ 6.4 51.3 $\pm$ 4.7 85.0 $\pm$ 16.1 62.0 $\pm$ 9.5
LinOSS-IMEX 75.5 $\pm$ 4.3 87.5 $\pm$ 4.0 58.9 $\pm$ 8.1 29.9 $\pm$ 1.0 57.9 $\pm$ 5.3 80.0 $\pm$ 2.7 65.0 $\pm$ 8.5
LinOSS-IM 75.8 $\pm$ 3.7 87.8 $\pm$ 2.6 58.2 $\pm$ 6.9 29.9 $\pm$ 0.6 60.0 $\pm$ 7.5 95.0 $\pm$ 4.4 67.8 $\pm$ 9.7
Transformer 70.5 $\pm$ 0.1 84.3 $\pm$ 6.3 49.1 $\pm$ 2.5 40.5 $\pm$ 6.3 50.5 $\pm$ 3.0 OOM 59.0 $\pm$ 8.0
RFormer 72.5 $\pm$ 0.1 81.2 $\pm$ 2.8 52.3 $\pm$ 3.7 34.7 $\pm$ 4.1 55.8 $\pm$ 6.6 90.3 $\pm$ 0.1 64.5 $\pm$ 8.4
LrcSSM (Ours) 72.7 $\pm$ 5.7 85.2 $\pm$ 2.1 53.9 $\pm$ 7.2 36.9 $\pm$ 5.3 58.6 $\pm$ 3.1 90.6 $\pm$ 1.4 66.3 $\pm$ 8.3
@@
In the rightmost column of Table 1, we also report the average accuracy across all six datasets considered from the UEA-MTSCA archive. LrcSSM achieved an accuracy of $66.3\%$, placing it at the forefront alongside the LinOSS-IM model, outperforming all other state-of-the-art models, including LRU, S5, S6, Mamba, and LinOSS-IMEX. The implicit integration scheme of the LinOSS-IM model, seems to have played an important role, and we plan to investigate a similar integration scheme for LrcSSM, too. Our current scheme is just a simple explicit Euler.
5.2 Generalizing the Technique of Diagonal Model Design
Other non-linear RNNs can be transformed into an efficient, parallelizable form by following steps similar to those outlined in Section 3. In Appendix D, we provide a sketch of how this transformation of any non-linear recurrent model can be achieved, and in Table 2, we present results comparing LrcSSM with other non-linear SSM models constructed using the same diagonal model design. In the line of this work, we extend our method to build up other non-linear SSMs from Minimal Gated Units (MGU) [45], Gated Recurrent Units (GRU) [2], and Long Short-Term Memory (LSTM) [16], which we refer to as MguSSM, GruSSM and LstmSSM, respectively. Our experiments show that the LrcSSM model outperforms the other non-linear models across the six classification benchmarks.
Table 2: Experimentation with different non-linear RNN models formulated into efficient SSMs. For these experiments, we use a fixed configuration with input encoding of 64 and 6 blocks of SSMs with 64 units. For each dataset, the average test accuracy and standard deviation (%) are reported, and the last row presents the mean and standard error aggregated across datasets.
MguSSM GruSSM LstmSSM LrcSSM (ours)
Heart $74.0\pm 4.8$ $\mathbf{75.7\pm 4.7}$ $\mathbf{75.0\pm 3.5}$ $\mathbf{75.0\pm 2.6}$
SCP1 $78.3\pm 6.6$ $80.2\pm 4.2$ $78.8\pm 3.1$ $\mathbf{84.8\pm 2.8}$
[…]
Ethanol $31.1\pm 4.2$ $34.5\pm 1.9$ $32.6\pm 5.6$ $\mathbf{36.1\pm 1.1}$
[…]
Worms $\mathbf{90.0\pm 5.2}$ $86.1\pm 6.3$ $82.2\pm 4.5$ $85.6\pm 5.4$
Average $63.2\pm 8.8$ $63.1\pm 8.4$ $62.34\pm 8.0$ $\mathbf{65.4\pm 8.0}$
5.3 Ablation Studies
In Appendix E, we present ablation studies to analyze the contributions of different design choices. First, we compare liquid capacitance (as in LRCs (6)) with constant capacitance (as in STCs (5)) when scaled up to SSMs. Second, we show that simplifying the model to enforce a diagonal Jacobian does not degrade performance, based on experiments against the original full LRC model with dense Jacobians. Finally, we evaluate the impact of input- and state-dependency, as well as the use of real versus complex-valued matrices.
@@ 6 Discussion
Our experimental evaluations show that the LrcSSM model performs moderately well on short-horizon datasets, while demonstrating highly competitive performance on datasets with long input sequences, as shown in Table 1. In those long-sequence tasks, LrcSSMs outperform LRUs, Mamba, and S6, and also achieve better average performance across all datasets.
The only model that LrcSSMs generally does not outperform is the LinOSS-IM model, except on the EthanolConcentration dataset (for both LinOSS-IMEX and LinOSS-IM versions), and on MotorImagery and EigenWorms (in the case of LinOSS-IMEX). This may be attributed to the fact that LinOSS is based on forced linear second-order ODEs, whereas LrcSSMs are built upon LRCs, which are non-linear first-order ODEs. Another possible reason lies in the integration technique: while we were able to outperform the implicit-explicit (IMEX) integration scheme, we did not surpass the fully implicit one (IM) in average test accuracy. This suggests that more sophisticated integration schemes for LrcSSMs (which currently use explicit Euler) may be worth investigating.
@@
In this paper, we focused on the biologically inspired non-linear LRC model, and demonstrated how this model can be made more efficient for long-sequence modeling, by redesigning its underlying state-recurrence matrix $\mathbf{A}$ and its input-transition vector $\mathbf{b}$, such that the resulting Jacobian is a diagonal matrix, for the state-update iterations. This matrix can then be directly used in the parallelizable methods of [8], which gives an exact parallel update, and not an approximation. As we have shown, this approach can also be applied to other non-linear RNNs of interest.
@@
As pointed out in Section A.2, this parallelized version holds a good promise towards efficient non-linear RNNs compared to sequential computation costs. Linear SSMs have also the same costs. However, we also have to take into account that LRCs need multiple Newton steps to converge at each iteration, which linear SSMs do not require. The number of iterations depends on the convergence of the state updates, which stops once the difference between the consecutive state updates gets below a defined threshold (see Line 4 of Algorithm 1). The number of iterations per datasets for LrcSSM is shown in Figure 2.